Please help me write two equations for this word problem. One equation for members and one for non members.
To encourage customers a movie theater is offering memberships. The membership costs $75.00 per year plus $2.00 per movie. For non-members, the cost of a movie is $5.75.
Thanks sooo much! Lisa

Answer by Theo(13342)

in the following equations:

y = total cost
m = cost per movie
x = number of movies per year
f = fixed cost per year

general equation is:

y = m*x + f

for members:

m = $2.00 per movie
f = $75.00 per year

member equation is:

y = 2*x + 75

for non-members:

m = $5.75

non-member equation is:

y = 5.75*x

break even point is when the 2 equations are equal to each other.

2*x + 75 = 5.75*x

subtract 2*x from both sides of the equation to get:

3.75*x = 75

x = 75/3.75 = 20

At 20 movies in a year, they break even. above 20 movies in a year, the total cost to the member becomes less than the total cost to the non-member.

member equation is:

y = 2x + 75
when x = 20, y = 40+75 = 115

non-member equation is:

y = 5.75*x
when x = 20, y = 20*5.75 = 115

anything over 20, the member pays less.

assume x = 30

member pays 2*30 + 75 = 60 + 75 = 135
non-member pays 5.75*30 = 172.5